OpenStudy (anonymous):

graphing form : parallel to 3x-6y=5 , containing (-2,4)

OpenStudy (aroub):

-6y=-3x+5 y=1/2x-5/6 parallel=same slope y=mx+b , m= 1/2 , (-2,4) 4=1/2(-2)+b 4=-1+b 5=b y=1/2x +5